OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike? The red bike is $____. (Be sure to type only the price of the red bike. DO NOT use a dollar sign ($) in your answer.)

OpenStudy (ganpat):

x = price for red y = price for blue.. x + y = 300 and y = x -14

OpenStudy (ganpat):

x + x -14 = 300 2x = 286.. x= 143 y = 300- 143 = 157..

OpenStudy (anonymous):

x=157